# Aperiodic sets of three types of convex polygons

### v1

@@ 1 Introduction
The “ASP comprising three convex polygons (= the set of three convex polygons that are aperiodic with no matching condition on the edges)” in the above problem is known for the two convex pentagons and one convex hexagon created by Ammann, as shown in Figure 1, based on the Penrose tiles111 The Penrose tiles are considered to be an ASP comprising two convex polygons (rhombuses) with matching condition on the edges [1, 3, 12]. [3, 12]. In the tiling of the two convex pentagons and one convex hexagon shown in Figure 1, the reflected (flipped) convex pentagons and hexagon were not used. The tiling is edge-to-edge222 A tiling by convex polygons is edge-to-edge if any two convex polygons in a tiling are either disjoint or share one vertex or an entire edge in common. Then other case is non-edge-to-edge [3, 11, 12].
@@
Let us consider “ASP comprising $n$ types of polygons ($n$ is an integer greater than or equal to one)” by focusing on “a set of three convex polygons that are aperiodic” in the above-mentioned problem. For the polygons (tiles) included in the ASP and the set of original problems, if the polygons are congruent, they are treated as the same type; however, from the text (or symbols) alone, it is not clear whether the anterior or posterior side polygons are treated as the same type. Further, similar to the problem of monohedral tiling, it is possible that polygons in those cases may sometimes be treated as the same type of tiles in the reflected image.
@@
Following this policy, the case in [9] is “ASP comprising two types of concave polygons,” and the case in [10] is “ASP comprising one type of concave diagram444 Tile(1,1)11(1,1)( 1 , 1 ) corresponding to $a=b=1$ of Tile$(a,b)$ shown in [9] is a concave polygon and a concave diagram, and Spectres in [10] are concave diagrams. From a viewpoint that does not distinguish between anterior and posterior sides, both cases of [9] and [10] are considered “ASP comprising one type of concave diagram.””. Moreover, the case shown by Ammann in Figure 1 is “ASP comprising three types of convex polygons (two types of convex pentagons and one type of convex hexagon).”
@@
Note that if we extract the convex polygons of $n-1$ or fewer types included in the set, we assume that they cannot admit a tiling. It is because it is natural that the convex polygons of $n-1$ or fewer types extracted from the set cannot generate periodic tiling where $n\geq 2$; however, if they can generate non-periodic tiling, it is replaced by the result that there exists an ASP comprising fewer types of convex polygons.
@@ 2 Dividing Tile(1,1)11(1,1)( 1 , 1 ) into ASP comprising convex polygons
Tile(1,1)11(1,1)( 1 , 1 ), corresponding to $a=b=1$ in Tile$(a,b)$ shown in [9], can generate a periodic tiling if the use of reflected tiles is allowed during the tiling generation process (see Figure 2(b)). However, as shown in [10], Tile(1,1)11(1,1)( 1 , 1 ) can only generate non-periodic tiling if and only if it does not allow the use of reflected tiles (see Figure 3). Using this property, ASP comprising three types of convex polygons can be obtained from Tile(1,1)11(1,1)( 1 , 1 ). This method divides the interior of Tile(1,1)11(1,1)( 1 , 1 ) into five convex polygons with five or more edges666 Because all triangles and quadrilaterals are monotiles (which do not require the use of reflected tiles to form periodic tilings) the number of edges of convex polygons contained in ASP is at least five. Then, there are no convex polygonal monotiles with seven or more edges [3, 5]., and we have so far identified four methods (Methods 1–4). The tile sets created by the division of the four methods all contain the convex pentagon $P_{1}$, which cannot generate a tiling (that is, no monotile) in the upper-left corner of Figure 4 ($P_{1}$ is convex pentagons with line symmetry and cannot distinguish between the anterior and posterior sides). Methods 1 (Figure 4), 2 (Figure 5), and 4 (Figure 18) divide Tile(1,1)11(1,1)( 1 , 1 ) to create five convex polygons containing three $P_{1}$. In Method 3 (Figure 15), one $P_{1}$ (= $CP(X_{3})$) is included in the five convex polygons formed by dividing Tile(1,1)11(1,1)( 1 , 1 ). We note that the tilings generated by the ASP, comprising three types of convex polygons created by Methods 1–4, are non-edge-to-edge.
@@ 2.1 Method 1 and Method 2
  • In the case of $\beta=90^{\circ}$ in Method 2 (see Figure 5(d)), there is the convex pentagon monotile $P_{2}$ belonging to both the Type 2 and Type 4 families with the property of line symmetry that cannot distinguish between the anterior and posterior sides (see Appendix A); therefore, the set of three types of convex polygons by this division is not an ASP.
  • In the cases of $\alpha=105^{\circ}$ in Method 1 (see Figure 4(e)) and $\beta=105^{\circ}$ in Method 2 (see Figure 5(e)), the three convex polygons created by division are identical. The set of two convex polygons created by this division is not an ASP, because the convex hexagon created by this division is a convex hexagonal monotile belonging to the Type 1 family, which can form periodic tiling without using reflected tiles (see Appendix B).
@@
[…]
@@
From the cases “(a) $75^{\circ}<\alpha<105^{\circ}$, (b) $105^{\circ}<\alpha<165^{\circ}$, (c) $165^{\circ}<\alpha<180^{\circ}$” of Method 1 shown in Figure 4 and “(a) $90^{\circ}<\beta<105^{\circ}$, (b) $105^{\circ}<\beta<165^{\circ}$, (c) $165^{\circ}<\beta<180^{\circ}$” of Method 2 shown in Figure 5, candidates of ASP comprising three types of convex polygons can be created. We refer to it as “candidate” because if we generate tilings with convex polygons contained in the tile set, we have to make sure that there exist only non-periodic tilings using the shape of Tile(1,1)11(1,1)( 1 , 1 ), if and only if all types of convex polygons contained in the tile set are used. For example, the division of $\alpha=90^{\circ}$ in “(a) $75^{\circ}<\alpha<105^{\circ}$” of Method 1 creates a convex pentagon monotile $P_{2}$ belonging to both the Type 2 and Type 4 families with line symmetry, as shown in Figure 8, which can form periodic tiling (see Appendix A). Thus, the set of three types of convex polygons created by the division of $\alpha=90^{\circ}$ in Method 1 is not an ASP.
As concrete examples of the “candidate” cases, we investigated the sets of three types of convex polygons in the case of $\alpha=98^{\circ}$ in Method 1 shown in Figure 9 and in the case of $\alpha=124^{\circ}$ in Method 1 shown in Figure 10. We confirmed that these tile sets were ASPs comprising three types of convex polygons. Figures 11 and  12 show the non-periodic tilings generated by the ASP comprising three types of convex polygons in the cases of $\alpha=98^{\circ}$ and $\alpha=124^{\circ}$ in Method 1.
@@ 2.3 Method 4
The fourth division method is Method 4, as shown in Figure 18. This can be regarded as an improved version of the division shown in Figure 4(e). Based on Figure 4(e), one of the four $P_{1}$s was transformed into a division containing a convex pentagonal monotile belonging only to the Type 13 family.
Figure. 18: Case in Method 4.
It can be considered that the tile set of $P_{1}$, $P_{2}$, and $P_{3}$ in Figure 18 is not an ASP, because the set contains the convex pentagonal monotile $P_{3}$ belonging to the Type 13 family. However, in this study, we assumed that the reflected tiles cannot be used during all tiling generation processes. $P_{3}$, which belongs to the Type 13 family, requires the use of reflected tiles during the tiling generation process (see Figure 18)777 If convex pentagonal monotiles do not belong to the Types 1 and 3–5 families, then the convex pentagonal monotiles cannot generate tilings without using reflected tiles (see Appendix A). The convex pentagonal monotiles belonging to the Type 13 family cannot belong to the Types 1 and 3–5 families at the same time due to the tile conditions (i.e., the convex pentagonal monotiles belonging to the Type 13 families and those belonging to the Types 1 and 3–5 families are disjoint) [3, 11, 12, 18].. The tile set of the three types of convex pentagons created by this division does not include reflected $P_{3}$. In other words, tiling cannot be generated with only this $P_{3}$ under the condition that “reflected tiles cannot be used during all tiling generation processes” (we actually confirmed this fact).
We examined whether tiling is possible using $P_{1}$, $P_{2}$, and $P_{3}$, as shown in Figure 18, and confirmed that the only combination that can construct a structure in which the tiling does not collapse using $P_{1}$, $P_{2}$, and $P_{3}$ is that of dividing Tile(1,1)11(1,1)( 1 , 1 ) (using the shape of Tile(1,1)11(1,1)( 1 , 1 )), as shown in Figure 18. Therefore, under the condition that the tiles in the tile set cannot be reflected (the use of reflected tiles is not allowed) during the tiling generation process, the tile set of the three types of convex pentagons that can be created using Method 4 in Figure 18 is an ASP. Figure 19 shows the non-periodic tiling generated by the ASP comprising three types of convex pentagons in Figure 18.
Figure. 19: Non-periodic tiling generated by the ASP comprising three types of convex pentagons in Figure 18.
@@ 3 Conclusion
Because the ASP comprising three types of convex polygons shown in this study is based on Tile(1,1)11(1,1)( 1 , 1 ) shown in [9] and assumes that Tile(1,1)11(1,1)( 1 , 1 ) is the chiral aperiodic monotile shown in [10], it is not confirmed that these sets are actually ASP. Thus, because [10] is still a preview version, it does not hold if Tile(1,1)11(1,1)( 1 , 1 ) is not a chiral aperiodic monotile (thus far, there is no evidence for such a case).
If the types and number of tiles in a tile set can be freely increased under the condition that the tiles in the tile set cannot be reflected during the tiling generation process, several ASPs comprising multiple types of convex polygons can be created. There will be various methods for dividing Tile(1,1)11(1,1)( 1 , 1 ) into four, five, or more types of convex polygons that are likely to be ASP (see Appendix D). This is because the convex pentagon can be divided into six types. Thus, by such a division, it would be easy to create convex pentagons that are not monotiles, or a convex pentagon containing only one side of a convex pentagonal monotile that requires the use of reflected tiles during the tiling generation process, as in Method 4. However, even if the tile sets that increase the number of tiles in this manner are aperiodic, they are not considered interesting. Therefore, we did not discuss the effects of increasing the type or number of such tiles in this study.
Acknowledgments. I would like to express my deep gratitude to Professor Emeritus Yoshio Agaoka, Hiroshima University, for his helpful comments and advice during the preparation of this manuscript. I also received constructive comments from Professor Shigeki Akiyama, University of Tsukuba, and Yoshiaki Araki.
@@
Currently, there are 15 known families of convex pentagonal monotiles called “Type,” as shown in Figure 20 [2, 3, 4, 6, 7, 11, 12, 18]. The tilings depicted with each convex pentagon in Figure 20 are representative of this Type. A representative tiling can be formed using only the relationships that can be derived from the tile conditions of each Type. As shown in Figure 20, the representative tilings of the Types 2 and 6–15 families use reflected tiles (convex pentagons). In other words, if convex pentagonal monotiles do not belong to the Types 1 and 3–5 families, then the convex pentagonal monotiles cannot generate tilings without using reflected tiles.
@@
Figure 22 shows the periodic tiling formed by convex pentagonal monotiles belonging to both the Type 2 and Type 4 families with line symmetry in the cases of $\beta=90^{\circ}$ in Method 2 (Figure 5(d)) and $\alpha=90^{\circ}$ in Method 1 (Figure 8).
Figure. 20: Fifteen Type families of convex pentagonal monotiles. Each of the convex pentagonal monotiles is defined by some conditions between the lengths of the edges and the magnitudes of the angles, but some degrees of freedom remain. For example, a convex pentagonal monotile belonging to Type 1 family satisfies that the sum of three consecutive angles is equal to $360^{\circ}$. This condition for Type 1 family is expressed as $A+B+C=360^{\circ}$ in this figure. Only the pentagonal monotile of Types 14 and 15 families have one degree of freedom, except for size. For example, the value of $C$ of the pentagonal monotile of Type 14 family is $\cos^{-1}((3\sqrt{57}-17)/16)\approx 1.2099\;\mbox{rad}\approx 69.32^{\circ}$. The gray region in each tiling indicates a translation unit (a unit that can generate a periodic tiling by translation only).
@@
Figures 24(a) and 24(b) show the periodic tilings formed by convex hexagonal monotiles belonging to the Type 1 family in the cases of $\alpha=75^{\circ}$ (Figure 4 (d)) and $\alpha=105^{\circ}$ (Figure 4(e)) in Method 1, respectively.
@@
Figure. 25: Periodic tiling with $P_{1}$, $P_{2}$ and reflected $P_{2}$ in the case of $\alpha=180^{\circ}$ in Method 1 (see Figure 4(g)). The reflected $P_{2}$ are given an asterisk mark “*.”
Appendix Appendix D ASP comprising four or five types of convex polygons888 The contents (the results of study) in this appendix were obtained prior to the main results of this manuscript.

### v3

@@ 1 Introduction
The “ASP comprising three convex polygons (= the set of three convex polygons that are aperiodic with no matching condition on the edges)” in the above problem is known for the two convex pentagons and one convex hexagon created by Ammann, as shown in Figure 1, based on the Penrose tiles111 The Penrose tiles are considered to be an ASP comprising two convex polygons (rhombuses) with matching condition on the edges [1, 3, 12]. [3, 12]. In the tiling of the two convex pentagons and one convex hexagon shown in Figure 1, the reflected (flipped) convex pentagons and hexagon were not used. The tiling is edge-to-edge222 A tiling by convex polygons is edge-to-edge if any two convex polygons in a tiling are either disjoint or share one vertex or an entire edge in common. Then other case is non-edge-to-edge [3, 11, 12]..
@@
Let us consider “ASP comprising $n$ types of polygons ($n$ is an integer greater than or equal to one)” by focusing on “a set of three convex polygons that are aperiodic” in the above-mentioned problem. For the polygons (tiles) contained in the ASP and the set of original problems, if the polygons are congruent, they are treated as the same type; however, from the text (or symbols) alone, it is not clear whether the anterior or posterior side polygons are treated as the same type. Further, similar to the problem of monohedral tiling, it is possible that polygons in those cases may sometimes be treated as the same type of tiles in the reflected image.
@@
Following this policy, the case in [9] is “ASP comprising two types of concave polygons,” and the case in [10] is “ASP comprising one type of concave diagram444 Tile(1,1)11(1,1)( 1 , 1 ) corresponding to $a=b=1$ of Tile$(a,b)$ shown in [9] is a concave polygon and a concave diagram, and Spectres in [10] are concave diagrams. From a viewpoint that does not distinguish between anterior and posterior sides, both cases of [9] and [10] are considered “ASP comprising one type of concave diagram.” Note that Tile(1,1)11(1,1)( 1 , 1 ) with 13-edges is considered an equilateral 14-edges polygon.”. Moreover, the case shown by Ammann in Figure 1 is “ASP comprising three types of convex polygons (two types of convex pentagons and one type of convex hexagon).”
@@
Note that if we extract the convex polygons of $n-1$ or fewer types contained in the set, we assume that they cannot admit a tiling. It is because it is natural that the convex polygons of $n-1$ or fewer types extracted from the set cannot generate periodic tiling where $n\geq 2$; however, if they can generate non-periodic tiling, it is replaced by the result that there exists an ASP comprising fewer types of convex polygons.
@@ 2 Dividing Tile(1,1)11(1,1)( 1 , 1 ) into ASP comprising convex polygons
Tile(1,1)11(1,1)( 1 , 1 ), corresponding to $a=b=1$ in Tile$(a,b)$ shown in [9], can generate a periodic tiling if the use of reflected tiles is allowed during the tiling generation process (see Figure 2). However, as shown in [10], Tile(1,1)11(1,1)( 1 , 1 ) can only generate non-periodic tiling if and only if it does not allow the use of reflected tiles (see Figure 3). Using this property, ASP comprising three types of convex polygons can be obtained from Tile(1,1)11(1,1)( 1 , 1 ). This method divides the interior of Tile(1,1)11(1,1)( 1 , 1 ) into five convex polygons with five or more edges666 Because all triangles and quadrilaterals are monotiles (which do not require the use of reflected tiles to form periodic tilings) the number of edges of convex polygons contained in ASP is at least five. Then, there are no convex polygonal monotiles with seven or more edges [3, 5]., and we have so far identified four methods (Methods 1–4). The tile sets created by the division of the four methods all contain the convex pentagon $P_{1}$, which cannot generate a tiling (that is, no monotile) in the upper-left corner of Figure 4 ($P_{1}$ is convex pentagons with line symmetry and cannot distinguish between the anterior and posterior sides). Methods 1 (Figure 4), 2 (Figure 5), and 4 (Figure 18) divide Tile(1,1)11(1,1)( 1 , 1 ) to create five convex polygons containing three $P_{1}$. In Method 3 (Figure 15), one $P_{1}$ (= $CP_{1}(X_{3})$) is contained in the five convex polygons formed by dividing Tile(1,1)11(1,1)( 1 , 1 ). We note that the tilings generated by the ASP, comprising three types of convex polygons created by Methods 1–4, are non-edge-to-edge.
@@ 2.1 Method 1 and Method 2
  • In the case of $\beta=90^{\circ}$ in Method 2 (see Figure 5(d)), there is the convex pentagonal monotile $P_{2}$ belonging to both the Type 2 and Type 4 families with the property of line symmetry that cannot distinguish between the anterior and posterior sides (see Appendix A); therefore, the set of three types of convex polygons by this division is not an ASP.
  • In the cases of $\alpha=105^{\circ}$ in Method 1 (see Figure 4(e)) and $\beta=105^{\circ}$ in Method 2 (see Figure 5(e)), the convex polygons created by division are identical. The set of two types of convex polygons created by this division is not an ASP, because the convex hexagon $P_{3}$ created by this division is a convex hexagonal monotile belonging to the Type 1 family, which can form periodic tiling without using reflected tiles (see Appendix B).
@@
Figure. 6: Periodic tiling with $P_{1}$ and $P_{2}$ with $\alpha=165^{\circ}$ in Method 1 ($\beta=165^{\circ}$ in Method 2).
@@
From the cases “(a) $75^{\circ}<\alpha<105^{\circ}$, (b) $105^{\circ}<\alpha<165^{\circ}$, (c) $165^{\circ}<\alpha<180^{\circ}$” of Method 1 shown in Figure 4 and “(a) $90^{\circ}<\beta<105^{\circ}$, (b) $105^{\circ}<\beta<165^{\circ}$, (c) $165^{\circ}<\beta<180^{\circ}$” of Method 2 shown in Figure 5, candidates of ASP comprising three types of convex polygons can be created. We refer to it as “candidate” because if we generate tilings with convex polygons contained in the tile set, we have to make sure that there exist only non-periodic tilings using the shape of Tile(1,1)11(1,1)( 1 , 1 ), if and only if all types of convex polygons contained in the tile set are used. For example, the division of $\alpha=90^{\circ}$ in “(a) $75^{\circ}<\alpha<105^{\circ}$” of Method 1 creates a convex pentagonal monotile $P_{2}$ belonging to both the Type 2 and Type 4 families with line symmetry, as shown in Figure 8, which can form periodic tiling (see Appendix A). Thus, the set of three types of convex polygons created by the division of $\alpha=90^{\circ}$ in Method 1 is not an ASP.
As concrete examples of the “candidate” cases, we investigated the sets of three types of convex polygons in the case of $\alpha=98^{\circ}$ in Method 1 shown in Figure 9 and in the case of $\alpha=124^{\circ}$ in Method 1 shown in Figure 10. We confirmed that these tile sets are ASPs comprising three types of convex polygons. Figures 11 and  12 show the non-periodic tilings generated by the ASP comprising three types of convex polygons in the cases of $\alpha=98^{\circ}$ and $\alpha=124^{\circ}$ in Method 1.
@@ 2.3 Method 4
The fourth division method is Method 4, as shown in Figure 18. The results of the tile sets (five convex polygons containing three $P_{1}$) created by the division of Method 4 differ depending on the value of the parameter $\gamma$ shown in Figure 18.
[…]
The cases “(e) $\gamma=150^{\circ}$, (f) $\gamma=165^{\circ}$” of Method 4 shown in Figure 18 are not ASPs. The reasons are shown below.
  • In the case of $\gamma=150^{\circ}$ in Method 4 (see Figure 18(e)), there is the convex hexagonal monotile $P_{3}$ belonging to the Type 1 family that can form periodic tiling without using reflected tiles (see Appendix B); therefore, the set of two types of convex polygons by this division is not ASP. Note that this division is the same as for $\alpha=105^{\circ}$ in Method 1 (see Figure 4(e)).
  • In the case of $\gamma=165^{\circ}$ in Method 4 (see Figure 18(f)), there is the convex pentagonal monotile $P_{2}$ belonging to both the Type 2 and Type 4 families with the property of line symmetry that cannot distinguish between the anterior and posterior sides (see Appendix A); therefore, the set of three types of convex polygons by this division is not an ASP.
[…]
Figure. 18: Cases in Method 4.
It can be considered that the tile set of $P_{1}$, $P_{2}$, and $P_{3}$ in the case of $\gamma=75^{\circ}$ in Method 4 (see Figure 19) is not an ASP, because the set contains the convex pentagonal monotile $P_{2}$ belonging to the Type 13 family. However, in this study, we assumed that the reflected tiles cannot be used during all tiling generation processes. $P_{2}$, which belongs to the Type 13 family, requires the use of reflected tiles during the tiling generation process (see Figure 19)888 If convex pentagonal monotiles do not belong to the Types 1 or 3–5 families, then the convex pentagonal monotiles cannot generate tilings without using reflected tiles (see Appendix A). The convex pentagonal monotiles belonging to the Type 13 family cannot belong to the Types 1 or 3–5 families at the same time due to the tile conditions (i.e., the convex pentagonal monotiles belonging to the Type 13 families and those belonging to the Types 1 or 3–5 families are disjoint) [3, 11, 12, 18].. The tile set of the three types of convex pentagons created by this division does not contain reflected $P_{2}$. In other words, tiling cannot be generated with only this $P_{2}$ under the condition that “reflected tiles cannot be used during all tiling generation processes” (we actually confirmed this fact). We examined whether tiling is possible using $P_{1}$, $P_{2}$, and $P_{3}$, as shown in Figure 19, and confirmed that the only combination that can construct a structure in which the tiling does not collapse using $P_{1}$, $P_{2}$, and $P_{3}$ is that of dividing Tile(1,1)11(1,1)( 1 , 1 ) (using the shape of Tile(1,1)11(1,1)( 1 , 1 )), as shown in Figure 19. Therefore, under the condition that the tiles in the tile set cannot be reflected during the tiling generation process, the tile set of the three types of convex pentagons that can be created using the case of $\gamma=75^{\circ}$ in Method 4 is an ASP.
Figure. 19: Case of $\gamma=75^{\circ}$ in Method 4.
Figure. 20: Non-periodic tiling generated by the ASP comprising three types of convex pentagons in in the case of $\gamma=75^{\circ}$ in Method 4.
[…]
Figure. 22: Non-periodic tiling generated by the ASP comprising three types of convex pentagons in in the case of $\gamma=180^{\circ}$ in Method 4.
@@ 3 Conclusion
Because the ASPs comprising three types of convex polygons shown in this study are based on Tile(1,1)11(1,1)( 1 , 1 ) shown in [9] and assumes that Tile(1,1)11(1,1)( 1 , 1 ) is the chiral aperiodic monotile shown in [10], it is not confirmed that these sets are actually ASPs. Thus, because [10] is still a preview version, it does not hold if Tile(1,1)11(1,1)( 1 , 1 ) is not a chiral aperiodic monotile (thus far, there is no evidence for such a case).
If the types and number of tiles in a tile set can be freely increased under the condition that the tiles in the tile set cannot be reflected during the tiling generation process, several ASPs comprising multiple types of convex polygons can be created. There will be various methods for dividing Tile(1,1)11(1,1)( 1 , 1 ) into four, five, or more types of convex polygons that are likely to be ASP (see Appendix D). This is because the convex pentagon can be divided into six types. Thus, by such a division, it would be easy to create convex pentagons that are not monotiles, or a convex pentagon containing only one side of a convex pentagonal monotile that requires the use of reflected tiles during the tiling generation process, as in the case of $\gamma=75^{\circ}$ in Method 4. However, even if the tile sets that increase the number of tiles in this manner are aperiodic, they are not considered interesting. Therefore, we did not discuss the effects of increasing the type or number of such tiles in this study.
Acknowledgments. The authors would like to express my deep gratitude to Professor Emeritus Yoshio Agaoka, Hiroshima University, for his helpful comments and advice during the preparation of this manuscript. The authors also received constructive comments from Professor Shigeki Akiyama, University of Tsukuba, and Yoshiaki Araki.
@@
Currently, there are 15 known families of convex pentagonal monotiles called “Type,” as shown in Figure 23 [2, 3, 4, 6, 7, 11, 12, 18]. The tilings depicted with each convex pentagon in Figure 23 are representative of this Type. A representative tiling can be formed using only the relationships that can be derived from the tile conditions of each Type. As shown in Figure 23, the representative tilings of the Types 2 and 6–15 families use reflected tiles (convex pentagons). In other words, if convex pentagonal monotiles do not belong to the Types 1 or 3–5 families, then the convex pentagonal monotiles cannot generate tilings without using reflected tiles.
@@
Figure 25 shows the periodic tiling formed by convex pentagonal monotiles belonging to both the Type 2 and Type 4 families with line symmetry in the cases of $\beta=90^{\circ}$ in Method 2 (Figure 5(d)), $\alpha=90^{\circ}$ in Method 1 (Figure 8), and $\gamma=165^{\circ}$ in Method 4 (Figure 18(f)).
Figure. 23: Fifteen Type families of convex pentagonal monotiles. Each of the convex pentagonal monotiles is defined by some conditions between the lengths of the edges and the magnitudes of the angles, but some degrees of freedom remain. For example, a convex pentagonal monotile belonging to Type 1 family satisfies that the sum of three consecutive angles is equal to $360^{\circ}$. This condition for Type 1 family is expressed as $A+B+C=360^{\circ}$ in this figure. The pentagonal monotiles of Types 14 and 15 families have one degree of freedom, that of size. For example, the value of $C$ of the pentagonal monotile of Type 14 family is $\cos^{-1}((3\sqrt{57}-17)/16)\approx 1.2099\;\mbox{rad}\approx 69.32^{\circ}$. The gray region in each tiling indicates a translation unit (a unit that can generate a periodic tiling by translation only).
@@
Figures 27(a) and 27(b) show the periodic tilings formed by convex hexagonal monotiles belonging to the Type 1 family in the cases of $\alpha=75^{\circ}$ (Figure 4 (d)) and $\alpha=105^{\circ}$ (Figure 4(e)) in Method 1, respectively.
@@
Figure. 28: Periodic tiling with $P_{1}$, $P_{2}$, and reflected $P_{2}$ in the case of $\alpha=180^{\circ}$ in Method 1 (see Figure 4(g)). The reflected $P_{2}$ are given an asterisk mark “*.”
Appendix Appendix D ASPs comprising four or five types of convex polygons999 The contents (the results of study) in this appendix were obtained prior to the main results of this manuscript.